### v1

# AutoTAMP: Autoregressive Task and Motion Planning with LLMs as Translators and Checkers

Abstract
For effective human-robot interaction, robots need to understand, plan, and execute complex, long-horizon tasks described by natural language. The recent and remarkable advances in large language models (LLMs) have shown promise for translating natural language into robot action sequences for complex tasks. However, many existing approaches either translate the natural language directly into robot trajectories, or factor the inference process by decomposing language into task sub-goals, then relying on a motion planner to execute each sub-goal. When complex environmental and temporal constraints are involved, inference over planning tasks must be performed jointly with motion plans using traditional task-and-motion planning (TAMP) algorithms, making such factorization untenable. Rather than using LLMs to directly plan task sub-goals, we instead perform few-shot translation from natural language task descriptions to an intermediary task representation that can then be consumed by a TAMP algorithm to jointly solve the task and motion plan. To improve translation, we automatically detect and correct both syntactic and semantic errors via autoregressive re-prompting, resulting in significant improvements in task completion. We show that our approach outperforms several methods using LLMs as planners in complex task domains. [^1] Datasets and Codes are available at https://github.com/yongchao98/AutoTAMP
Keywords: Large Language Models, Task and Motion Planning, Human-Robot Interaction
1 Introduction
A long-standing goal of robotics has been to provide agents with the ability to find and execute optimal plans for complex tasks. Robots not only need to reason about the task in the environment and find a satisfying sequence of actions but also verify the feasibility of executing those actions according to the robot’s motion capabilities. This problem is referred to as task and motion planning (TAMP), and there has been considerable research on efficient algorithms [1]. Classic solutions rely on specifying tasks in a dedicated planning representation amenable to such algorithms [2, 3, 4, 5].
While this approach to task specification has been quite successful, directly using these representations requires training and experience, making them a poor candidate as an interface for non-expert users. As an alternative, natural language (NL) provides an intuitive and flexible mode of task descriptions. Excitingly, pre-trained large language models (LLMs) have demonstrated surprisingly good performance on many language-related tasks [6], and there has been an associated burst of research investigating their various potential applicability to TAMP [7, 8, 9, 10, 11, 12, 13].
Early efforts showed promising results using LLMs as direct task planners [7], i.e., generating a sequence of sub-tasks based on a set of natural language instructions, but were limited by a lack of feedback and verification of executability of the sub-task sequence. Further research addressed executability by connecting sub-tasks to control policy affordance functions [8], providing environmental feedback of robot actions [10], and interleaving action feasibility checking with LLM action proposals [9]. However, these approaches struggle with various task complexities, such as temporally-dependent multi-step actions, action sequence optimization, task constraints, etc. Furthermore, these frameworks factor the planning problem and use LLMs to infer a task plan separately from the motion plan. In many situations, the task and motion plan must optimized together to fulfill the task. For instance, when the task is ‘reach all locations via the shortest path’, the order of places to be visited (task planning) depends on the geometry of the environment and the related motion optimization. Unfortunately, as we show in this paper, LLMs do not seem capable of directly generating trajectories, possibly related to limitations in complex spatial and numerical reasoning [14, 15].
Notably, the task representations used in classic TAMP, such as PDDL [2] or Temporal Logics [3], are sufficiently expressive to specify task complexities from which existing planning algorithms can find and verify action sequences satisfying those specifications. To benefit from both the user-friendliness of NL and the capabilities of existing TAMP algorithms, we approach the problem by using LLMs to translate from high-level task descriptions to formal task specifications. We are not the first to use LLMs as translators of task specifications [16, 17], but our work contributes several new ideas that improve the performance of such approaches. First, previous work focused on translating natural language into PDDL goals [16] or Linear Temporal Logics (LTL) [18], which only considered the problem of task planning. Here we utilize Signal Temporal Logic (STL) as the intermediary representation, from which the planner directly optimizes the whole trajectory, i.e, task and motion planning together, thus improving planning success rate. Second, we also use re-prompting to automatically correct task specification errors, both via an existing syntactic error correction framework and a novel semantic error checking and correcting loop using LLMs, resulting in significantly improved performance. We also conduct an ablation study integrating a fine-tuned NL-to-STL model [19] with our AutoTAMP approach and show that GPT-4 few-shot learning is competitive with fine-tuning. We conclude that in-context learning with pre-trained LLMs is well suited for language-to-task-specification translation for solving TAMP problems.
In this paper, we contribute the following:
  • Pre-trained LLM (GPT-3 and GPT-4) few-shot translation from NL task descriptions to STLs, and execution of trajectories via an STL planner.
  • A novel autoregressive re-prompting technique for automatic checking and correcting of semantic errors in generated task specifications via LLMs. We show that, when combined with automatic syntactic correction, this technique greatly improves task success rates.
  • Comprehensive experiments in challenging task domains, including several multi-agent tasks. We show that our approach outperforms direct LLM planning methods for tasks with hard geometric and temporal constraints.
  • In addition to our code, we publish a dataset of 1400 test cases consisting of the language instruction, environments, produced STL specifications, and resulting trajectories.
2 Problem Description
Figure 1: GPT-4 failure case for direct end-to-end trajectory planning. The orange line shows the correct path obeying the instruction. The purple and gray dashed lines show the trajectories from GPT-4 after first and second prompts, respectively. GPT-4 generates a list of $(x,y)$ locations with associated timepoints. The initial prompt describes the language modeling task, environment state, and instruction. Each object is a rectangle described by $(x,y)$ boundaries.
As shown in Figure 1, we aim to convert a natural language instruction, including spatial and temporal constraints, into a motion plan for a robot encoded as a set of timed waypoints, e.g., $(x_{i},y_{i},t_{i})$. The environment state is encoded as set of named obstacles described as polygons and is provided as additional context. Our task is to generate a constraint-satisfying trajectory based on the given instruction and the environment state. The robot must not surpass its maximum velocity, and the total operation time should not exceed the task time limit. We assume that the full trajectory is a linear interpolation between the timed waypoints; complex trajectories can be specified by dense waypoint sequences.
3 Methods
Figure 2: Illustration of different approaches applying LLMs for task and motion planning.
Figure 2 illustrates three of the approaches we compare in our work, each using LLMs in some capacity. Each takes as input (1) a text-based representation of the global environment state, (2) in-context examples for few-shot learning, and (3) a natural language instruction. For methods using feedback or re-prompting, we describe details of the re-prompt context in Appendix A.
3.1 LLM End-to-end Motion Planning
One natural idea is to use an LLM to handle both the task and motion planning by directly generating a trajectory for a given language instruction; we refer to this as LLM End-to-end Motion Planning. Where the generated trajectory violates constraints, we re-prompt the model with the constraint violation to produce another trajectory, allowing up to five such re-prompts. Figure 1 illustrates this pipeline, including a specific failure case with two constraint-violating trajectories.
3.2 LLM Task Planning
A more common approach is to use an LLM to handle the task planning by directly generating a sequence of sub-tasks from a given language instruction; we refer to this as LLM Task Planning. To generate a final trajectory, the sub-tasks are handled by an independent motion planner. In this work, these sub-tasks are limited to navigation actions, and the motion planning is handled by the STL planner used by our proposed approach; this permits fair comparison of results across methods. Each sub-task is converted to STL to be consumed by the planner (more details are presented in Appendix B). We evaluate and compare against three methods that each use LLMs for task planning: (1) Naive Task Planning, (2) SayCan, and (3) LLM Task Planning + Feedback.
Naive Task Planning As proposed by Huang et al. [7], we evaluate using LLMs to generate the entire sub-task sequence without checking for executability.
SayCan Alternatively, an LLM can be autoregressively prompted to generate each subsequent sub-task conditioned on the previous sub-tasks in the sequence. The next sub-task can be selected from the top K candidates by combining the language model likelihood with a feasibility likelihood of the candidate action and choosing the most-likely next sub-task. This is the method proposed by Ahn et al. [8]. We set K to 5 in our evaluations.
LLM Task Planning + Feedback A third task planning method combines full sequence generation with feasibility checking to both find sub-task sequences that satisfy the full task and verify their feasibility before execution. For any infeasible sub-tasks, the LLM can be re-prompted with feedback about the infeasible actions to generate a new sub-task sequence. This is similar to the hierarchical method proposed by Lin et al. [9] but with feedback for re-prompting.
Figure 3: An example of full prompt used by our AutoTAMP framework
3.3 Autoregressive LLM Specification Translation&Checking + Formal Planner
Unlike LLM Task Planning, our approach translates NL to STL with an LLM and then plans the trajectory with an STL planner, as shown in Figure 2. We include two re-prompting techniques to improve translation performance: one for syntactic errors and another for semantic errors. We show one example of the prompt in Figure 3.
Signal Temporal Logic Syntax In this work, we use STL [20] as a formal task specification that supports continuous real-time constraints suitable for time-critical missions. An STL formula is defined recursively according to the following syntax:
$\phi::=\pi^{\mu}\;|\;\neg\phi\;|\;\phi\land\varphi\;|\;\phi\lor\varphi\;|\;\textbf{F}_{[a,b]}\phi\;|\;\textbf{G}_{[a,b]}\phi\;|\;\phi\textbf{U}_{[a,b]}\varphi$ (1)
where $\phi$ and $\varphi$ are STL formulas, and $\pi^{\mu}$ is an atomic predicate. $\neg$ (negation), $\land$ (and), $\lor$ (or), $\Rightarrow$ (imply), and $\Leftrightarrow$ (equal)) are logical operators. $\textbf{F}_{[a,b]}$ (eventually/finally), $\textbf{G}_{[a,b]}$ (always/globally), and $\textbf{U}_{[a,b]}$ (until) are temporal operators with real-time constraints $t\in[a,b]$. Temporal operators with time constraints are illustrated by Appendix C, and other operators can be presented using the basic syntax. The action primitives in this work are ’enter(room_name)’ and ’not_enter(room_name)’. To avoid issues of parentheses matching, the generated STLs adhere to pre-order form, as discussed in Appendix D.
STL Trajectory Planner We use a state-of-the-art STL planner [21] capable of multi-agent planning. It defines the validity of an STL formula with respect to a trajectory and then optimizes the trajectory to maximize the validity. The planner only considers the state within or out of a region, and outputs a sequence of location & time pairs.
Syntactic Checking & Semantic Checking Open-loop translation can suffer from syntactic and semantic errors. We use two re-prompting techniques to automatically correct such errors. Like Skreta et al. [22] we use a verifier to check for syntax errors; any found are provided as feedback when re-prompting the LLM to generate corrected STL. We repeat until no errors are found (up to five iterations). For semantic errors, we propose a novel autoregressive re-prompting technique that feeds the STL planner’s generated state sequence (i.e., [[in(road), 0], [in(red kitchen), 0.5], [in(blue restroom2), 1.2]…]) back into the LLM to check whether it satisfies the original instruction. If it does not, the LLM is prompted to modify the STL, which repeats the syntactic and semantic re-prompting. This process terminates in the case of no detected errors or no change in STL (up to three iterations).
4 Experimental Design
Each of our task scenarios is set in a 2D environment and entails navigation of one or more robots; the robots have extent in the environment and are initialized with varying start positions. Each environment is comprised of regions with a shape, location, and characteristics (e.g., color, name, function). For each method we evaluate, the LLM is initially prompted with a description of the language task (e.g. task planning) and five in-context examples (these depend on the language task). To mitigate variance across prompts, we initially tested six different sets of examples for each method and chose the one that performed best. Through this testing, we found that the variance over prompts was insignificant relative to overall performance.
We evaluated the different methods described in Section 3 across six different task scenarios (three single-agent and three multi-agent) with different combinations of geometric and temporal constraints. For each scenario description below, we indicate the presence of these constraints below with G and T, respectively. For each method, we evaluate performance with both GPT-3 and GPT-4 as the LLM. Note that in multi-agent scenarios, we do not test SayCan or LLM Task Planning + Feedback because these methods are not straight-forwardly adaptable for multiple agents. We also terminate and report failure for test cases that take more than 90 minutes. We automatically check resulting trajectories via hard-coded checkers. The full set of experiments took two weeks using four 16-core CPUs; the cost of LLM API calls was ~1500USD.
Figure 4: HouseWorld and Chip’s Challenge are single-agent scenarios. Overcooked, Rover, and Wall are multi-agent scenarios. The black square in Overcooked is inadmissible. The lines indicate the correct trajectories following the instructions. For the HouseWorld and Chip’s Challenge environments, the black round dot and pentagonal dot indicate the start and end positions, respectively.
HouseWorld1 (single-agent) As shown in Figure 4(a), this is a house environment from Finucane et al. [23]. We first manually constructed 10 different instructions of varying complexity (see Appendix E) before prompting GPT-4 to paraphrase each into 9 differently worded instructions with the same meaning, resulting in 100 total instructions for this environment. For each instruction, we randomly initialize between two start-end position pairs for 200 total test cases. For this scenario, we do not impose a hard time constraint for the planned trajectory.
Example instruction: ’Visit two rooms with color closest to red, but not the pure red color.’
Corresponding STL: (finally room_purple and finally room_pink).
HouseWorld2 (T, single-agent) This scenario is identical to HouseWorld1, but each planned trajectory is subjected to a hard time constraint. This time limit is pre-determined by completing the correct trajectory with 0.8 maximum velocity.
Example instruction: ’Visit two rooms with color closest to red, but not the pure red color. The task should be completed within 10 seconds.’
Corresponding STL: (finally[0, 10] room_purple and finally[0, 10] room_pink).
The remaining task scenarios were designed with specific rules and goals for the agent(s) to follow. For each scenario, GPT-4 was used to paraphrase the original description into 20 uniquely worded variants with the same meaning. We instantiate three different instances of the environment for each scenario and randomize five different start/end location pairs for a total of 300 test cases. Note that the avoidance of collision among multiple agents has been inherently encoded in the STL planner.
Chip’s Challenge (G, single-agent) Figure 4(b) shows a scenario inspired by a level from Chip’s Challenge, a classic puzzle solving game with strict geometric and logical constraints. The robot must reach all goal regions (blue) but must acquire a unique key to pass through the corresponding door. The task planning must obey both the spatial and logical constraints.
Example instruction: ’Try to reach all the goals but you have to reach the corresponding key first to open the specific door. For example, you have to reach key1 ahead to open door1. Also remember always do not touch the walls.’
[…]
Overcooked (G & T, multi-agent) Figure 4(c) shows a scenario inspired by Overcooked, a popular cooking simulation game with strict time constraints. The agents must cooperatively gather ingredients and return to CookingRoom in a limited time. In this scenario, the multi-agent motion planning is challenged by limited space for agents to maneuver.
Example instruction: ’Enter all the ingredient room to pick up food. Once entered ingredient rooms, go to cooking room within 3 seconds. And all the agents should not collide each other and black obstacles.’
[…]
Rover (G & T, multi-agent) Figure 4(d) is a scenario used by Sun et al. [21]. Multiple agents must reach each observation region (blue goals) before transmitting the observations by entering a red region, all while subjected to time and energy constraints.
Example instruction: ’All rovers must reach the blue charging station within 5 units of time each time they exit it. Once they reach their destination, they need to get to a yellow transmitter within 2 time units to send the collected information to the remote control. Rovers must keep clear of black walls and other rovers. All target areas need to be visited.’
[…]
Wall (G & T, multi-agent) Figure 4(e) is also a scenario from Sun et al. [21]. Multiple agents must occupy each goal region (blue) while subject to a time constraint and a maneuver bottleneck.
Example instruction: ’Reach all the goals and stay there. Take care of the collision among each agent and to the walls.’
[…]
Appendix F shows key frame sequences of correctly planned trajectories for the three multi-agent scenarios as produced via our AutoTAMP method.
5 Results
5.1 Task Success Rates
Table 1: Task success rates in single-agent scenarios. The scenarios have hard temporal or geometric constraints.
HouseWorld1 HouseWorld2 Chip Challenge
(soft time cst.) (hard time cst.) (hard geo. cst.)
GPT-3 end2end $0\%$ $0\%$ $0\%$
GPT-3 naive task planning $74\%$ $36\%$ $0\%$
SayCan $75.5\%$ $36\%$ $0\%$
GPT-3 task planning/feed. 79% $40\%$ $0\%$
GPT-3/STL $28\%$ $27\%$ $29\%$
GPT-3/STL/Syn. $49\%$ $47\%$ $66\%$
GPT-3/STL/Syn./Sem. (AutoTAMP) $62\%$ 62% 74.3%
GPT-4 end2end $9.5\%$ $9.5\%$ $0\%$
GPT-4 naive task planning $90\%$ $45\%$ $0\%$
SayCan $90\%$ $47.5\%$ $0\%$
GPT-4 task planning/feed. 92% $49\%$ $0\%$
GPT-4/STL $43.5\%$ $42\%$ $42.7\%$
GPT-4/STL/Syn. $59.5\%$ $59\%$ $70\%$
GPT-4/STL/Syn./Sem. (AutoTAMP) $82.5\%$ 82% 87.7%
Table 2: Task success rates in multi-agent scenarios. Each scenario has hard temporal and geometric constraints.
Overcooked Rover Wall
(multi-agent, hard time & geo. cst.)
GPT-3 end2end $0\%$ $0\%$ $0\%$
GPT-3 naive task planning $13.3\%$ $0\%$ $7\%$
GPT-3/STL $25\%$ $22\%$ $74\%$
GPT-3/STL/Syn. $70\%$ $35\%$ $85\%$
GPT-3/STL/Syn./Sem. (AutoTAMP) 89% 60.7% 89.7%
GPT-4 end2end $5\%$ $0\%$ $6\%$
GPT-4 naive task planning $17\%$ $0\%$ $47\%$
GPT-4/STL $85\%$ $46\%$ $95\%$
GPT-4/STL/Syn. $94\%$ $67\%$ $95\%$
GPT-4/STL/Syn./Sem. (AutoTAMP) 100% 79% 100%
We report the task success rates for the single-agent and multi-agent scenarios in Table 1 and Table 2, respectively. For HouseWorld1 (Figure 4(a)) with no hard time constraint, we find that all methods using LLMs as task planners outperform our approach; this environment permits direct trajectories between any two positions and thus lacks geometric challenges that such methods will struggle with. When adding a strict time constraint (HouseWorld2), we see that such methods perform much worse while AutoTAMP’s success rate persists. For the remaining tasks that include geometric constraints, LLM End-to-end Motion Planning and Naive Task Planning both perform quite poorly. Unsurprisingly, we observe a general trend that GPT-4 outperforms GPT-3 for all methods.
5.2 Failure examples for LLM Task Planning and AutoTAMP Methods
LLM Task Planning: We observe that LLM Task Planning methods may fail at tasks that require long-horizon planning. Figure 5 below illustrates two examples. In (a), these methods inefficiently grab the key for a particular door only when that door is the next immediate obstacle. Rather than grab both key1 and key2 that are near the starting position before unlocking door1, it sequences the actions to grab key1, unlock door1, then grab key2. This inefficiency can lead to a violation of time constraints that are part of the task specification, resulting in task failure. In (b), ingredient1 is unnecessarily (and inefficiently) visited by both agents.
Figure 5: Two examples of failures by LLM Task Planning methods.
AutoTAMP: AutoTAMP (and the ablations) primarily fail because of translation errors. While the re-prompting techniques we use to improve translation are effective, there are still translation errors that result in task failure.
5.3 Planning Time
We report timing information for AutoTAMP with GPT-4 in Table 3. For each scenario, we report 10th, 50th, and 90th percentile runtime cost for the following steps in the overall process: GPT-4 API call, STL planner planning time, code execution time, GPT-4 translation to STL, syntactic check loop, semantic check loop. Overall, the STL planner runtime is the main bottleneck. In cases where the GPT-4 API call takes an excessive amount of time (i.e. more than 100s), we note that this is likely due to unstable latency and may not be a good proxy for decoding runtime.
We also report the number of iterations for both the syntactic and semantic re-prompting techniques for each task scenario. Note that semantic re-prompting subsequently requires an additional round of syntactic re-prompting for the new STL output; thus, syntactic re-prompting can have more total iterations than the limit of 5 since this limit is applied per STL output.
Table 3: Timing information for each step of AutoTAMP using GPT-4; Iteration counts for the syntactic and semantic re-prompting. For each, We report the 10th, 50th, and 90th percentile.
Time (s) HouseWorld2 Chip. Overcooked Rover Wall
Total Time $15/42/221$ $403/871/1579$ $34/77/157$ $79/123/174$ $21/43/67$
GPT-4 API $2/3/55$ $8/21/160$ $3/5/7$ $10/14/18$ $4/6/7$
STL Planner $10/35/152$ $310/732/1482$ $30/74/148$ $68/89/142$ $17/34/49$
Code Processing $/3/3$ 3 $3/4/4$ $/3/3$ 3 $3/4/4$ $/3/3$ 3
GPT-4/STL $5/9/10$ $172/457/681$ $16/20/45$ $15/18/29$ $6/9/15$
Syntactic Check $1/3/60$ $11/23/163$ $4/7/12$ $5/6/7$ $2/3/5$
Semantic Check $6/36/157$ $141/630/850$ $17/58/106$ $75/102/134$ $14/31/54$
Iterations HouseWorld2 Chip. Overcooked Rover Wall
Syntactic Check $1/2/2$ $3/5/7$ $1/2/2$ $2/3/6$ $1/2/3$
Semantic Check $1/2/3$ $/3/3$ 3 $2/2/3$ $2/3/3$ $2/2/3$
5.4 Ablation Studies
Syntactic and Semantic Checkers In Table 1 and 2, we also evaluate the impact of syntactic and semantic error correction on using LLMs to translate to STL. The results show that translation with no error correction has modest success across task scenarios, but both syntactic and semantic error correction significantly improve performance; this trend is present across all scenarios.
Fine-tuned NL2TL Translation Model Prior work (NL2TL [19]) has investigated translating from natural language to STL via a modular pipeline that utilizes a fine-tuned LLM. While incorporating NL2TL into our AutoTAMP framework was not part of our primary investigation, we report here its impact on task success in Table 4. The top three rows report AutoTAMP + NL2TL with ablations over syntactic and semantic re-prompting. The bottom three rows report the original ablation results of AutoTAMP without NL2TL. We see that integrating NL2TL generally provides a modest improvement. However, it is exciting that AutoTAMP appears competitive without the use of a fine-tuned model since it does not rely on a targeted dataset or additional offline training.
Table 4: Ablation studies of the AutoTAMP with/without integrating the finetuned NL2TL model.
HouseWorld2 Chip. Overcooked Rover Wall
GPT-4/STL/NL2TL $53\%$ $61\%$ $87\%$ $58\%$ $95\%$
GPT-4/STL/Syn./NL2TL $64\%$ $73\%$ $94\%$ $70\%$ $95\%$
GPT-4/STL/Syn./Sem./NL2TL 83.5% 86% 100% 79.7% 100%
GPT-4/STL $42\%$ $42.7\%$ $85\%$ $46\%$ $95\%$
GPT-4/STL/Syn. $59\%$ $70\%$ $94\%$ $67\%$ $95\%$
GPT-4/STL/Syn./Sem. 82% 87.7% 100% 79% 100%
6 Related Work
LLMs for TAMP Recent claims about the impressive reasoning capabilities of LLMs [6, 24] have led to interest in such models for task and motion planning. One approach is to directly use LLMs as planners [7, 8, 9, 10, 11, 12, 13, 25]. Initial work showed that naive zero-shot generation of a primitive action sequence from a high-level task description had relatively poor executability, but few-shot in-context learning, constraining output to admissible actions, and autoregressive step-by-step action generation significantly improved performance [7]. Subsequent efforts connected to motion planning by grounding the primitive actions to motion control policies and using affordance functions to guide LLM-based task and motion planning [8, 9]; Huang et al. [10] closed the loop by prompting with feedback, such as action success indication, before querying for the next action. Other work focused on how prompting can inform task execution, such as few-shot summarization for user task preferences [11] or extraction of commonsense arrangements for underspecified object placement [13]. Despite these successes, however, there is some evidence that LLMs exhibit poor performance on more realistic tasks [15, 26], motivating different approaches. While we are interested in LLMs for TAMP, our work does not directly use LLMs as planners.
Translating Language to Task Representations A natural alternative is to rely on dedicated planners by mapping from natural language to a planning representation. For over a decade, research has investigated mapping language to lambda calculus [27], motion planning constraints [28], linear temporal logic [29, 30, 31], and signal temporal logic [32, 33], among others. To address challenges of data availability, task generalization, and linguistic complexity, recent work has applied LLMs to this translation problem. Modular approaches have used LLMs to extract referring expressions with corresponding logic propositions to then construct a full temporal logic specification [34, 19]. Relying on LLMs for direct translation, other work has mapped from language to PDDL goals [17] or full PDDL problems [35, 16]. Our work similarly translates to a task specification, but we can represent complex constraints (e.g. temporal), and we introduce a novel mechanism for automatic detection and correction of semantic errors.
Re-prompting of LLMs The quality of LLM output is greatly improved with useful context, such as few-shot in-context learning for novel tasks [6]. LLMs for TAMP are typically also provided task-relevant information, such as environment state or admissible actions [12]. Re-prompting with additional context based on LLM output has been shown to be extremely beneficial, such as with iterative action generation [7], environmental feedback [10], inadmissible actions [7, 8, 9], unmet action preconditions [36, 35], code execution errors [37], and syntactic errors in structured output [22]. Our work utilizes the same syntactic correction re-prompting technique as Skreta et al. [22], but we further introduce automatic detection and correction of semantic errors via re-prompting.
7 Conclusion and Limitations
This paper presented AutoTAMP, a framework for using pre-trained LLMs as translators from language task descriptions to formal task specifications in STL via few-shot learning; a key contribution is that we show that our use of STL significantly improves the performance of the planner, and that the translation via automatic correction of syntax errors and a novel autoregressive re-prompting technique for semantic error checking and correction lead to further performance gains. Our experimental results show using LLMs to translate to task specifications that can be solved via a formal planner outperforms approaches that use LLMs directly as planners when handling tasks with complex geometric and temporal constraints.
Limitations 1) Though our results rely on using the best prompt out of several candidates, the prompt used may not be the optimal for eliciting the best performance of our LLMs. As a result, the individual results for each method may be improvable, but we expect the trends to persist and will continue to support the conclusion that LLMs are not well suited for direct complex TAMP. 2) Though the task success rates of AutoTAMP are good, the cost of the planning time is high, especially for multiple syntactic and semantic re-prompting iterations. Further work is needed that focuses on how to accelerate formal planners and decrease LLM inference latency. 3) Some tasks still fail due to translation errors. As part of future work, we aim to explore whether feedback via a human-in-the-loop framework can both improve translation and reduce AutoTAMP iterations.
Acknowledgments
We thank OpenAI for approving the usage of GPT-3/4 APIs.
The authors would also like to thank Charles Dawson, Oswin So, and Yue Meng for the early-stage discussion and final reviews of the work.
This work was supported by ONR under Award N00014-22-1-2478 and MIT-IBM Watson AI Lab. However, this article solely reflects the opinions and conclusions of its authors.
Appendix A Recursive LLM Prompts
In this work, some methods need to re-prompt LLMs with context of the prompt history. Unlike the GPT-4 API that provides a template for this, the GPT-3 API template is for a single inference history; so, it is necessary to include any prompt history as context that indicates prior user input and LLM responses. We show an example in Figure 6.
Figure 6: Example of prompting GPT-3 API recursively. The prompt should indicate previous user prompts and LLM responses with tags (red).
Appendix B Convert Sub-task Sequence to STL
In order to evaluate LLM Task Planning methods, we use an STL planner [21] (this is the same planner we use for our AutoTAMP method) as the motion planner for generating trajectories. The LLM Task Planning methods generate output of the form [[$enter(room_{1})$, $t_{1}$], [$enter(room_{2})$, $t_{2}$],…, [$enter(room_{i})$, $t_{i}$], [$enter(room_{i+1})$, $t_{i+1}$]]. We automatically convert to STL to be consumed by the planner, i.e., finally[$t_{1}$, $t_{1}$] $enter(room_{1})$ and finally[$t_{2}$, $t_{2}$] $enter(room_{2})$ and finally[$t_{i}$, $t_{i}$] $enter(room_{i})$ and globally $not\_enter(walls)$. If the LLM Task Planning method also requires an executability check (e.g. SayCan, LLM Task Planing + Feedback), we apply the same conversion per sub-task; for example, given a sub-trajectory $i$, we would convert to STL of the form finally[$t_{i+1}-t_{i}$, $t_{i+1}-t_{i}$] $enter(room_{i+1})$ and globally $not\_enter(walls)$ to then be checked for executability.
Appendix C STL Syntax
Table 5: STL illustration
$\textbf{F}_{[a,b]}\phi$
True at time $t$ if there exists a time in the interval $[t+a,t+b]$ where $\phi$ is true.
$\phi\textbf{U}_{[a,b]}\varphi$
True at time $t$ if $\varphi$ is true for some time $t^{{}^{\prime}}$ in the interval $[t+a,t+b]$, and for all times between $t$ and $t^{{}^{\prime}}$, the formula $\phi$ holds.
$\textbf{G}_{[a,b]}\phi$
True at time $t$ if for all times in the interval $[t+a,t+b]$, the formula $\phi$ holds.
$\varphi\textbf{$\Rightarrow$}\textbf{F}_{[300,infinite]}\phi$
If $\varphi$ happens, then some time after the next 300 time steps $\phi$ should happen.
$\neg\varphi_{1}\textbf{$\lor$}(\varphi_{2}\textbf{U}_{[286,348]}\varphi_{3})$
$\varphi_{1}$ should not happen , or else $\varphi_{2}$ happens and continues until at some point during the 286 to 348 time units $\varphi_{3}$ happens .
Appendix D Pre-order and In-order STL Formats
As shown in Figure 7, each STL can be viewed as a binary tree with each node in the tree having at most two child nodes. When requesting NL-to-STL translation, the targeted tokens of STLs can be linearized in an in-order (left subtree, root, right subtree) or pre-order (root, left subtree, right subtree) manner. For this work, we prompt LLMs to generate STL in the pre-order format to avoid parentheses matching errors and for compatability with the STL planner. We note that the translation performance with GPT-4 is similar for both formats.
Figure 7: Illustration of pre-order and in-order formats of STL. (a) Different text representations of the formats for the same STL. (b) An example binary tree representation of STL.
Appendix E Manually Designed Instructions for HouseWorld Scenario
  1. 1.
    Navigate into the green bedroom and then go to the yellow restroom.
  2. 2.
    Travel to the yellow restroom, and then go to the pink bedroom. Visit both the kitchen and the dining room. Finally end in the red bathroom.
  3. 3.
    Navigate into the green bedroom and then go to the yellow restroom, but remember do not touch the livingroom at anytime.
  4. 4.
    Go to one room with cyan color, then enter the bedroom and stay there for 5 seconds, finally reach the restroom. remember always do not touch the two true blue areas.
  5. 5.
    Visit two rooms with color closest to red, but not the pure red color.
  6. 6.
    Visit every room that is not light pink or cyan.
  7. 7.
    Always go to the kitchen immediately before entering either bathroom. Go to the yellow room and the red room, but do not enter any blue rooms.
  8. 8.
    Move to the kitchen, then move to the bedroom without entering any blue rooms. Stay there for at least 10 seconds, then return to the kitchen via the living room.
  9. 9.
    Eventually end up in the master bedroom, but first go to both restrooms. After you have gone to each restroom, do not pass through any other rooms.
  10. 10.
    Visit every room at least once, but not go to the pink room until you are in a larger blue room.
Appendix F Keyframes of Trajectories for Multi-agent Scenarios.
Figure 8: Multi-agent task scenarios: Overcooked, Rover, and Wall. For Overcooked, the blue regions represent target ingredients, the center black square in an inadmissible region, and the yellow and green squares are agents. For Rover and Wall, the blue regions represent goals, the black lines represent walls, and the agents are the brown, yellow, orange, and green squares. In all scenarios, the agents occupy space and are consequently constrained by the inadmissible regions when moving through the environments.